# swvicad.bib

### v1

@@ Abstract
Magnonics is an emerging research field that addresses the use of spin waves (magnons), purely magnetic waves, for information transport and processing. Spin waves are a potential replacement for electric current in modern computational devices that would make them more compact and energy efficient. The field is yet little known, even among physicists. Additionally, with the development of new measuring techniques and computational physics, the obtained magnetic data becomes more complex, in some cases including 3D vector fields and time-resolution. This work presents an approach to the audio-visual representation of the spin waves and discusses its use as a tool for science communication exhibits and possible data analysis tool. The work also details an instance of such an exhibit presented at the annual international digital art exhibition Ars Electronica Festival in 2022.
@@ 1 Introduction
Spin-Wave Voices [^1] https://youtu.be/QGWoRnfgXfI?si=xlATu7Q-8XybePoM is an art installation, which was presented at a digital art exhibition Ars Electronica Festival“Welcome to Planet B” [^2] https://ars.electronica.art/planetb/en/spin-wave-voices/ . This exhibition allowed the wide public to explore the innovative technology of future computational devices: the spin waves. The processes behind the capabilities of modern digital devices happen on micro- and nanoscopic scales and at incredible speeds, making them impossible to observe with the naked eye and difficult to comprehend [Gelsinger2001, Nikonov2022]. Spin waves (magnons) are considered to be a potential replacement for electrons for data processing in novel computing devices [Chumak2022, Wang2018]. The Spin-Wave Voices exhibit allowed visitors to experience the future technology through visualization and auditory display (AD) in a playful, easy-going way. Furthermore, the people from the field were allowed to get insights into the data and spin-wave behavior.
Schematic of the installation is shown in Fig. 1. Every visitor had a chance to excite spin waves in 5 differently shaped microstrips (a) by pressing corresponding pedals (b), as described in more detail in Section  3. The waves were visualized and sonified in a way, that the spatial differences of the waves caused by the shapes of the microstrips were easily recognizable by visitors.
Figure 1: Pedal room schematic representation.
@@ 1.1 Spin Waves for Future Computing
One can approach this problem, e.g. optimizing computing strategies, as is already actively researched for AI, so-called AI research strategies [Verdecchia2023]. Another way would be making computing devices more energy efficient. Magnonics is a promising research field in this regard, as it employs spin waves (magnons), purely magnetic waves [Rana2019], for information transport and processing. Among other advantageous features, magnonic devices can be scaled down to atomic dimensions, can operate in a wide frequency range up to hundreds of THz, are able to process data at temperatures spanning from ultra-low to room temperature. Additionally, spin waves can make unconventional computing possible due to their nonlinear properties [Chumak2022]. Moreover, there has been increasing interest in studying spin waves, in various geometries, including curved structures, as 3D curvature can introduce unique magnetic properties and novel phenomena due to the interplay of curvature, topology, and magnetic interactions [Streubel2016, Fischer2020].
The research of magnonics, especially when it comes to the investigation of patterns in spin wave behavior, can often be impeded by an extensive complexity of the data, either measured or obtained from micromagnetic simulations [Donnelly2017]. Analysis of complex data requires the use of sophisticated visualization techniques in order to reduce the cognitive load on the analyst while keeping the visualization representative of the phenomenon [Crameri2020]. Spatial, especially 3D, data combined with an additional temporal dimension may make 3D representations and animations unavoidable in a comprehensive visualization.
@@ 1.2 Complex Data Visualization
Measurement or simulation data of physics phenomena, like spin waves, are too complex to understand in their raw format. Scientific visualization techniques are frequently applied to facilitate the understanding and interpretation of spatial (scalar) fields [schroeder_visualization_2006, telea_data_2014]. Such datasets include fixed spatial positions for the scalar values, leaving only a small subset of effective channels for encoding the scalar value itself [munzner2015visualization, p. 102]. Among the more effective channels for encoding a scalar value under these restrictions are depth (i.e., 3D position)—if the scalar field is a 2D plane—and color. The use of depth for encoding scalar fields results in a 3D visualization that comes with drawbacks like occlusion, distortion, or multiple viewpoints [munzner2015visualization, pp. 120-122]. Therefore, the use of 3D visualizations must be justified and the benefits must outweigh the drawbacks. One such benefit is to support people in building a mental model of 3D phenomena [munzner2015visualization, p. 124]. In particular, we can use waves on a water surface as a metaphor for spin waves to aid people in building a mental model about them. We can use alternative 3D representations (e.g., using slices of the data represented as a ridgelines plot [wilke_fundamentals_nodate]) to avoid occlusion.
@@ 1.3 Auditory Display and Sonification
The ability of AD to provide a better understanding, or an appreciation, of changes and structures in the sonified data has been attracting rapidly growing attention over the last years. One of the fields that benefited from the sonification and auditory display of complex data is astronomy [Zanella2022]. The reason for that is auditory display being one of the most appropriate approaches when the data has complex patterns and changes in time [Hermann2011].
With the development of measuring techniques [Donnelly2017, Weigand2022] and computational physics, e.g. micromagnetics [Abert2019, Leliaert2019], the field of magnonics meets similar challenges as in astronomy. The obtained data, which needs to be analyzed, becomes more complex, in some cases including 3D vector fields and time-resolution. Adding AD to visualization tools could benefit from human auditory perception, which allows for monitoring and processing multilayered auditory data. Additionally, the ability of humans to learn from sonified data and being able to notice that "something is wrong" [Hermann2011] can open new opportunities for getting unexpected insights from the data.
@@ 1.4 Spin-Wave Voices Objectives
In this work, an approach to combining data sonification with interactive visualization of spin waves, excited in confined Ni80Fe20 (Py) microstrips of different shapes, is presented. In particular, the influence of the shape of a structure on spin-wave patterns during a transient response (from the initial moment up to the stable state) of the system on uniform microwave (MW) field excitation is illustrated. Even though excitation conditions for all shapes are exactly the same, one can experience different spin-wave patterns depending on the shape. This can show one of many unique properties of spin waves, which make them a promising data carrier for future computational devices and trigger fruitful discussions leading to better awareness of the phenomena. A prototype developed following our approach was named Spin-Wave Voices and has been evaluated in the science-communication framework at a digital art exhibition Ars Electronica Festival 2022 "Welcome to Planet B" [^3] https://ars.electronica.art/planetb/en/spin-wave-voices/. The Spin-Wave Voices installation was aiming to:
@@
In the following sections detailed description of technical details of the project, its presentation and reception at the festival are discussed. Additionally, Spin-Wave Voices was presented at two research groups, who specialize on magnetic phenomena, which is also discussed in more detail in Section 4.2.
@@ 2.1 Spin-Wave Data Source
As a data source results of micromagnetic simulations of spin-wave transient response, meaning dynamics from the beginning of the excitation till the steady state is achieved, in Ni80Fe20 (Py) microstrips of various shapes are used. Spin waves in the structures are excited by applying an in-plane static magnetic field $\mu_{0}H=87$ mT in a direction indicated in Fig. 2, and an out-of-plane periodic magnetic field at excitation frequency $f_{\mathrm{MW}}=9.4$ GHz.
Micromagnetic simulations were performed using MuMax3 [mumax] with the following parameters [Pile2022]: frame size that included sample material and an empty space of $1.5\times 5~{}\mu m$, a thickness of 25 nm, saturation magnetization $M_{\mathrm{s}}=730$ kA/m, Gilbert damping parameter $\alpha=0.008$, exchange stiffness constant $A_{\mathrm{ex}}=13\times 10^{-12}$ J/m, and g-factor=2.12. As a result, 20 frames per excitation period of the spatially resolved magnetization configuration over the first 50 excitation periods were recorded from the simulations. As a result of the simulations for each shape, 1000 magnetization vector field configurations were obtained, depicting an evolution in time of the magnetization dynamics. Each configuration’s spatial distribution of the out-of-plane magnetization component was extracted as a 2D array of data (spin-wave profile). This resulted in 1000 data arrays $500\times 150$ for each shape, which were further used for the visualization and AD.
The shapes were chosen in a way that with the same excitation parameters, such as external static and excitation magnetic field, the spin-wave profiles would differ significantly. The resulting shapes are shown in Fig. 2. With the chosen parameters a uniform excitation is observed in the ellipse shape (b), and spin waves are excited in the other four shapes with different wave lengths.
@@ 2.2 Interactive Visualization
The interactive visualization for Spin-Wave Voices was created using Cables.gl [^4] https://cables.gl (cables). Cables is a free online tool for creating interactive visual web content. It was developed by the Berlin-based studio undev [^5] https://undev.studio initially as a development tool for interactive WebGL and Web Audio projects. It uses a node-based interface, allowing for quick prototyping.
@@
The goal of the visualization is to showcase the transient response of a micro-structure to uniform MW field excitation. This is achieved by leveraging a three-dimensional visualization, described above, combined with an animation. The generated visualization represents the transient response of a micro-structure by plotting the out-of-plane magnetization component on the outline of the structure at every moment in time during excitation. The plots are then used to create an animation. The out-of-plane component is encoded both with a diverging color scheme (color depending on the direction of the out-of-plane component) and as elevation/depression out of the plain of the structure outline, further highlighting the direction of magnetization.
As a result, the visualization appears as a surface: flat before the excitation (out-of-plain magnetization is zero) and colored, dynamically undulating with elevations and depressions demonstrating the transient behavior. The animation is controlled by the viewer: at any point, they are able to restart the visualization/excitation of the spin waves and observe the evolution of the process anew. To be perceivable by a human the process is visualized at a rate of half of the excitation period per second.
@@ 2.3 Sonification of Spin-Wave Data
Our simulation data is two-dimensional, time-dependent wave data and thus suitable for audification-type sonification [Hermann2011], with considerations given to the spatial and frequency characteristics of these measurements.
@@
Audio signals are sampled at a pre-defined audio sample rate, e.g. $44.1$ kHz from $500\times 150$ pixel image streams at a pre—defined image frame rate, e.g. $10$ fps. Periodic audio-rate sampling path $(x,y,t)$ through the spatio-temporal data grid illustrated in Fig. 3 (grey dashed arrow) is predefined. The period of the path defines the fundamental frequency of the resulting musical audio signal while the data itself (which can be also observed in the corresponding visualization) creates a dynamic, pulsating texture that gives each shape a certain timbre. To generate the audio data, the data is linearly interpolated at the locations defined by the sampling path. Because only one audio sample value per resampled image is calculated and to avoid wasting resources, only the path is resampled and not full images for each audio sample. The latter is achieved by linear interpolation from the image data "on the fly", by calculating one audio sample value for each $(x,y,t)$ value.
The simulation data is sonified from five different shapes, plus the installation setup resembles a piano-like instrument, therefore, the frequencies from a pentatonic scale (98 Hz, 110 Hz, 123 Hz, 165 Hz, 185 Hz) were chosen as fundamental frequencies that define the period of the sample paths.
The resulting audio data is written to files and imported into an Ableton Live’s Looper instrument. As the simulation data consists of a transient and a steady-state segment, the times of the segment boundaries are identified and loop points are set accordingly. When a pedal is pressed, the transient segment is played back once and the steady-state portion is repeated as long as the corresponding pedal is pressed.
As the floor pedals are set up from left to right, the corresponding loops are panned from left to right to reflect shapes’ spacial arrangement.
@@ 2.4 Hardware + Software
The physical realization of the installation consists of a screen, a projector, a sound system, five foot pedals that can be used by the visitors to trigger the visualization and the sonification (see Figs. 1 and 6), and a regular PC laptop for running the visualization and sonification processes (not show in the figures). As an input devices a standard sostenuto pedals connected to an Arduino microcontroller that translates the voltage changes to MIDI messages were used. The pedals were placed in a line on the floor, spaced widely enough to enable pressing no more than two pedals by a single visitor. These messages are then sent to the PC running Ableton Live for the sonification part, and a web browser running a cables.gl patch for the visualization.
@@ 3.1 Space A: Science Communication
  • The exhibit demonstrates the variety of spin-wave patterns occurring in confined microstrips of different shapes by means of an audio-visual display;
  • The possibility to change spin-wave patterns by just changing the shape of a wire demonstrates the flexibility of spin waves and, among other already mentioned advantages, their potential for the realization of non-conventional computing;
@@
To help make the introduction more engaging the room features a microscope allowing to demonstrate the scale of the microstrips by showing a rectangular strip [Fig. 2 (d)] next to a human hair. The room also has a scientific poster placed on the wall allowing more motivated visitors to delve deeper into the physics behind the spin waves. In addition, the walls are decorated with two cleanroom suits - a special kind of clothing worn in a cleanroom (space, which maintains a very low amount of airborne particulates) while producing microstrips. A cleanroom environment is required to prevent, e.g. dust particles, which are usually larger than the strips, from interfering with the production process. After the introduction, visitors are invited to experience the audio-visual display in Space B.

### v2

# refs2024.bib

@@ Abstract
Magnonics is an emerging research field that addresses the use of spin waves (magnons), purely magnetic waves, for information transport and processing. Spin waves are a potential replacement for electric current in novel computational devices that would make them more compact and energy efficient. The field is yet little known, even among physicists. Additionally, with the development of new measuring techniques and computational physics, the obtained magnetic data becomes more complex, in some cases including 3D vector fields and time-resolution. This work presents an approach to the audio-visual representation of the spin waves and discusses its use as a tool for science communication exhibits and possible data analysis tool. The work also details an instance of such an exhibit presented at the annual international digital art exhibition Ars Electronica Festival in 2022.
@@ 1 Introduction
Spin-Wave Voices is an art installation, which was presented at the digital art exhibition Ars Electronica Festival “Welcome to Planet B” [^1] https://ars.electronica.art/planetb/en/spin-wave-voices/ . This exhibition allowed the wide public to explore the innovative technology of future computational devices: the spin waves. The processes behind the capabilities of modern digital devices happen on micro- and nanoscopic scales and at incredible speeds, making them impossible to observe with the naked eye and difficult to comprehend [Gelsinger2001, Nikonov2022]. Spin waves (magnons) are considered to be a potential replacement for electrons for data processing in novel computational devices [Chumak2022, Wang2018]. The Spin-Wave Voices exhibit allowed visitors to experience the future technology through visualization and auditory display (AD) in a playful, easy-going way. Furthermore, the people from the field were allowed to get insights into the data and spin-wave behavior.
Schematic of the installation is shown in Fig. 1. Every visitor had a chance to excite spin waves in 5 differently shaped microstrips (a) by pressing corresponding pedals (b), as described in more detail in Section  3. The waves were visualized and sonified in a way, that the spatial differences of the waves caused by the variation of the shapes of the microstrips were easily recognizable by visitors.
Figure 1: Spin-Wave Voices’ Space B (see Section 3 and Fig. 6) schematic representation: (a) screen, (b) pedals, (c) loudspeakers, and (d) projector.
@@ 1.1 Spin Waves for Future Computing
One can approach this problem, e.g. optimizing computing strategies, as is already actively researched for AI, so-called AI research strategies [Verdecchia2023]. Another way would be making computing devices more energy efficient. Magnonics is a promising research field in this regard, as it employs spin waves (magnons), purely magnetic waves [Rana2019], for information transport and processing. Among other advantageous features, magnonic devices can be scaled down to atomic dimensions, can operate in a wide frequency range up to hundreds of THz, are able to process data at temperatures spanning from ultra-low to room temperature. Additionally, spin waves can transfer data without Joule heating [Chumak2014] and can make unconventional computing possible due to their nonlinear properties [Chumak2022]. Moreover, there has been increasing interest in studying spin waves, in various geometries, including curved structures, as 3D curvature can introduce unique magnetic properties and novel phenomena due to the interplay of curvature, topology, and magnetic interactions [Streubel2016, Fischer2020].
The research of magnonics, especially when it comes to the investigation of patterns in spin wave behavior, can often be impeded by an extensive complexity of the data, either measured or obtained from micromagnetic simulations [Donnelly2017]. Analysis of complex data requires the use of sophisticated visualization techniques in order to reduce the cognitive load on the analyst while keeping the visualization representative of the phenomenon [Crameri2020].
@@ 1.2 Complex Data Visualization
Measurement or simulation data of physics phenomena, like spin waves, are too complex to understand in their raw format. Scientific visualization techniques are frequently applied to facilitate the understanding and interpretation of spatial (scalar) fields [schroeder_visualization_2006, telea_data_2014]. Such datasets include fixed spatial positions for the scalar values, leaving only a small subset of effective channels for encoding the scalar value itself [munzner2015visualization, p. 102]. Among the more effective channels for encoding a scalar value under these restrictions are depth (i.e., 3D position) – if the scalar field is a 2D plane – and color. The use of depth for encoding scalar fields results in a 3D visualization that comes with drawbacks like occlusion, distortion, or multiple viewpoints [munzner2015visualization, pp. 120-122]. Therefore, the use of 3D visualizations must be justified and the benefits must outweigh the drawbacks. One such benefit is to support people in building a mental model of 3D phenomena [munzner2015visualization, p. 124]. In particular, we can use waves on a water surface as a metaphor for spin waves to aid people in building a mental model about them. We can use alternative 3D representations (e.g., using slices of the data represented as a ridgelines plot [wilke_fundamentals_nodate]) to avoid occlusion.
@@ 1.3 Auditory Display and Sonification
The ability of AD to provide a better understanding, or an appreciation, of changes and structures in the sonified data has been attracting rapidly growing attention over the last years. One of the fields that benefited from the sonification and auditory display of complex data is astronomy [Zanella2022]. The reason for that is auditory display being one of the most appropriate approaches when the data has complex patterns and changes in time [Hermann2011, chapter 2].
With the development of measuring techniques [Donnelly2017, Weigand2022] and computational physics, e.g. micromagnetics [Abert2019, Leliaert2019], the field of magnonics meets similar challenges as in astronomy. The obtained data, which needs to be analyzed, becomes more complex, in some cases including 3D vector fields and time-resolution. Adding AD to visualization tools could benefit from human auditory perception, which allows for monitoring and processing multilayered auditory data. Additionally, the ability of humans to learn from sonified data and being able to notice that “something is wrong” [Hermann2011, chapter 1] can open new opportunities for getting unexpected insights from the data.
@@ 1.4 Spin-Wave Voices Objectives
In this work, an approach to combining data sonification with interactive visualization of spin waves, excited in confined Ni80Fe20 (Py) microstrips of different shapes, is presented. In particular, the influence of the shape of a structure on spin-wave patterns during a transient response (from the initial moment up to the stable state) of the system on uniform microwave (MW) field excitation is illustrated. Even though excitation conditions for all shapes are exactly the same, one can experience different spin-wave patterns depending on the shape. This can show one of many unique properties of spin waves, which make them a promising data carrier for future computational devices and trigger fruitful discussions leading to better awareness of the phenomena. A prototype developed following our approach was named Spin-Wave Voices and has been evaluated in the science-communication framework at a digital art exhibition Ars Electronica Festival 2022 “Welcome to Planet B” [^2] https://ars.electronica.art/planetb/en/spin-wave-voices/. The Spin-Wave Voices installation was aiming to:
  1. (1)
@@
In the following sections detailed description of technical details of the project, its presentation and reception at the festival are discussed. Additionally, Spin-Wave Voices was presented at two research groups, who specialize on magnetic phenomena, and at the IEEE International Magnetics Conference 2024 (INTERMAG 2024) as a contributed talk, which is also discussed in more detail in Section 4.2.
@@ 2.1 Spin-Wave Data Source
As a data source results of micromagnetic simulations of spin-wave transient response (i.e. dynamics from the beginning of the excitation till the steady state is achieved) in Ni80Fe20 microstrips of various shapes are used. Spin waves in the structures are excited by applying an in-plane static magnetic field $\mu_{0}H=87$ mT in a direction indicated in Fig. 2, and an out-of-plane periodic magnetic field at excitation frequency $f_{\mathrm{MW}}=9.4$ GHz.
Micromagnetic simulations were performed using MuMax3 [mumax] with the following parameters [Pile2022]: frame size that included sample material and an empty space of $1.5\times 5~{}\mu m$, a thickness of 25 nm, saturation magnetization $M_{\mathrm{s}}=730$ kA/m, Gilbert damping parameter $\alpha=0.008$, exchange stiffness constant $A_{\mathrm{ex}}=13\times 10^{-12}$ J/m, and g-factor=2.12. As a result, 20 frames per excitation period of the spatially resolved magnetization configuration over the first 50 excitation periods were recorded from the simulations. As a result of the simulations for each shape, 1000 magnetization vector field configurations were obtained, depicting an evolution in time of the magnetization dynamics. Each configuration’s spatial distribution of the out-of-plane magnetization component was extracted as a 2D array of data (spin-wave pattern). This resulted in 1000 data arrays $500\times 150$ for each shape, which were further used for the visualization and AD.
The shapes were chosen in a way that with the same excitation parameters, such as external static and excitation periodic magnetic field, the spin-wave patterns would differ significantly. The resulting shapes are shown in Fig. 2. With the chosen parameters a uniform excitation is observed in the ellipse shape (b), and spin waves are excited in the other four shapes with different wave lengths.
@@ 2.2 Interactive Visualization
The interactive visualization for Spin-Wave Voices was created using Cables.gl [^3] https://cables.gl/p/WPjgr6 - for the Editor view only. (cables). Cables is a free online tool for creating interactive visual web content. It was developed by the Berlin-based studio undev [^4] https://undev.studio initially as a development tool for interactive WebGL and Web Audio projects. It uses a node-based interface, allowing for quick prototyping.
@@
The goal of the visualization is to showcase the transient response of a micro-structure to uniform MW field excitation. This is achieved by leveraging a three-dimensional visualization, described above, combined with an animation. The generated visualization represents the transient response of a micro-structure by plotting the out-of-plane magnetization component on the outline of the structure at every moment in time during excitation. The plots are then used to create an animation. The out-of-plane component is encoded both with a diverging color scheme, which indicates the direction of the out-of-plane component, and as elevation/depression out of the plain of the structure outline, indicating the magnitude of the magnetization.
As a result, the visualization appears as a surface: flat before the excitation (out-of-plain magnetization is zero) and, after excitation is applied colored, dynamically undulating with elevations and depressions demonstrating the transient behavior. The animation is controlled by the viewer: at any point, they are able to restart the visualization/excitation of the spin waves and observe the evolution of the process anew. To be perceivable by a human the process is visualized at a rate of half of the excitation period per second.
@@ 2.3 Sonification of Spin-Wave Data
The simulation data is two-dimensional, time-dependent wave data and thus suitable for audification-type sonification [Hermann2011, chapter 12], with considerations given to the spatial and frequency characteristics of these measurements.
@@
Audio signals are sampled at a predefined audio sample rate, e.g. $44.1$ kHz, from $500\times 150$ pixel image streams at a predefined image frame rate, e.g. $10$ fps. Periodic audio-rate sampling path $(x,y,t)$ through the spatio-temporal data grid illustrated in Fig. 3 with the gray dashed line with an arrow indicating the sampling direction is predefined. The sampling period of the path defines the fundamental frequency of the resulting musical audio signal while the data itself (which can be also observed in the corresponding visualization) creates a dynamic, pulsating texture that gives each shape a certain timbre. To generate the audio data, the data is linearly interpolated at the locations defined by the sampling path. Because only one audio sample value per resampled image is calculated and to avoid wasting resources, only the path is resampled and not full images for each audio sample. The latter is achieved by linear interpolation from the image data “on the fly”, by calculating one audio sample value for each $(x,y,t)$ value.
The simulation data is sonified from five different shapes, plus the installation setup resembles a piano-like instrument, therefore, the frequencies from a pentatonic scale (98 Hz, 110 Hz, 123 Hz, 165 Hz, 185 Hz) were chosen as fundamental frequencies that define the period of the sample paths. The resulting audio data is written to files and imported into an Ableton Live’s Looper instrument. As the simulation data consists of a transient and a steady-state segment, the times of the segment boundaries are identified and loop points are set accordingly. When a pedal is pressed, the transient segment is played back once and the steady-state portion is repeated as long as the corresponding pedal is pressed. As the floor pedals are set up from left to right, the corresponding loops are panned from left to right to reflect shapes’ spacial arrangement.
@@ 2.4 Hardware + Software
As schematically shown in Fig. 1 the physical realization of the installation consists of a screen (a), a projector (d), a sound system (c), five foot pedals (b) that can be used by the visitors to trigger the visualization and the sonification (see additionally Fig. 6), and a regular PC laptop for running the visualization and sonification processes (not shown in the figures). As an input devices a standard sostenuto pedals connected to an Arduino microcontroller that translates the voltage changes to MIDI messages were used. The pedals were placed in a line on the floor, spaced widely enough to enable pressing no more than two pedals by a single visitor. These messages are then sent to the PC running Ableton Live for the sonification part, and a web browser running a cables.gl patch for the visualization [^5] https://github.com/Spin-Wave-Voices/SWV-App - offline simplified version (no pedals)..
@@ 3.1 Space A: Science Communication
  • The possibility to change spin-wave patterns by just changing the shape of a wire, which is shown at the exhibit, demonstrates the flexibility of spin waves and, among other already mentioned advantages, their potential for the realization of non-conventional computing;
@@
To help make the introduction more engaging the room features a microscope allowing to demonstrate the scale of the microstrips by showing a rectangular strip [Fig. 2 (d)] next to a human hair (see Microscope image in Fig. 6). The room also has a scientific poster placed on the wall allowing more motivated visitors to delve deeper into the physics behind the spin waves. In addition, the walls are decorated with two cleanroom suits - a special kind of clothing worn in a cleanroom (space, which maintains a very low amount of airborne particulates) while producing microstrips. A cleanroom environment is required to prevent, e.g. dust particles, which are usually larger than the strips, from interfering with the production process. After the introduction, visitors are invited to experience the audio-visual display in Space B.